# line-of-sight UL uplink

### v1

@@ Abstract
We study the impact of imperfect line-of-sight (LoS) phase tracking on the performance of cell-free massive MIMO networks. Unlike prior works that assume perfectly known or completely unknown phases, we consider a realistic regime where LoS phases are estimated with residual uncertainty due to hardware impairments, mobility, and synchronization errors. To this end, we propose a Rician fading model where LoS components are rotated by imperfect phase estimates and attenuated by a deterministic phase-error penalty factor.
We derive a linear MMSE channel estimator that captures statistical phase errors and unifies prior results, reducing to the Bayesian MMSE estimator with perfect phase knowledge and to a zero-mean model in the absence of phase knowledge. To address the non-Gaussian setting, we introduce a virtual uplink model that preserves second-order statistics of channel estimation, enabling the derivation of tractable centralized and distributed MMSE beamformers. To ensure fair assessment of the network performance, we apply these beamformers to the true uplink model and compute the spectral efficiency bounds available in the literature.
@@ I Introduction
Existing studies typically analyze cell-free systems under two limiting cases of LoS phase knowledge. On one hand, studies such as [17, 2] and phase-aware case in [12, 16] assume perfect knowledge of the LoS phases, which enables tractable analysis but overlooks practical estimation errors. On the other hand, studies [16, 12] also consider the case of completely unknown LoS phases, thereby discarding valuable deterministic information inherent in the slowly varying LoS components. However, as also pointed out in [16, 12], the practical system performance fall in between these two scenarios. Furthermore, the performance evaluation of LoS-aware cell-free systems has largely been restricted to maximum-ratio (MR) combining [17, 16, 12]. While analytically convenient, MR combining does not fully exploit the beamforming potential of large-scale cell-free networks. Complementary studies [13, 5, 14] have investigated phase noise effects in Rayleigh fading channels, that is, without explicit modeling of the LoS components. These simplifications limit our understanding of the impact of partial LoS phase knowledge on channel estimation and beamforming in cell-free systems.
@@ II System and Channel Model
We consider a user-centric cell-free massive multiple-input multiple-output (mMIMO) network with $L$ access points, indexed by $\mathcal{L}:=\{1,\dots,L\}$, jointly serving $K$ users indexed by $\mathcal{K}:=\{1,\dots,K\}$, arbitrarily distributed across the service area. All APs are connected to a central processing unit (CPU) via an error-free fronthaul network [10]. Each AP is equipped with $N$ antennas, while each user has a single antenna.
@@ II-A Rician Channel under Imperfect Phase Tracking
In an arbitrary coherence block, let $\bm{h}_{k,l}\in\mathbb{C}^{N}$ denote a realization of the channel vector between user $k\in\mathcal{K}$ and the $N$ antennas of AP $l\in\mathcal{L}$. More specifically, the channel is modeled as $(\forall k\in\mathcal{K})(\forall l\in\mathcal{L})$
@@
where $\bm{\bar{h}}_{k,l}\in\mathbb{C}^{N}$ is the deterministic spatial signature of the LoS component, and $\mathbf{R}_{k,l}\in\mathbb{C}^{N\times N}$ is the spatial covariance matrix of the non-line-of-sight (NLoS) components $\bm{\tilde{h}}_{k,l}$, and $\theta_{k,l}\sim\mathcal{U}[-\pi,\pi]$ is the true phase of the LoS component. For each realization of $\theta_{k,l}$, an estimate
$\hat{\theta}_{k,l}=\theta_{k,l}-\varepsilon_{k,l}$
@@
However, as noted in [12, 16], practical systems operate between these extremes. In fact, the LoS phase often varies more slowly than the NLoS components, whose dynamics are dominated by delay and Doppler spreads. For example, recent work [17] claims that the LoS phase may remain approximately constant across multiple coherence blocks, particularly in the frequency domain. Therefore, with our model we focus on the intermediate case where the LoS phase $\theta_{k,l}$ varies more slowly than the NLoS components $\bm{\tilde{h}}_{k,l}$, but still fast enough such that it cannot be perfectly tracked. We denote by $\bm{\hat{\theta}}_{l}=[\hat{\theta}_{1,l},\dots,\hat{\theta}_{K,l}]^{\mathsf{T}}\in\mathbb{C}^{1\times K}$ and $\bm{\hat{\theta}}=[\bm{\hat{\theta}}_{1},\dots,\bm{\hat{\theta}}_{L}]\in\mathbb{C}^{L\times K}$ the phase estimates for all channels of AP $l$ and the network respectively.
The phase estimates $\bm{\hat{\theta}}$ are assumed to be available as common knowledge across the network. The phase estimation error $\varepsilon_{k,l}$ is modeled as uniformly distributed, $\varepsilon_{k,l}\sim\mathcal{U}[-\delta,\delta]$, with $0\leq\delta\leq\pi$, and it is assumed independent of the NLoS component. Under these assumptions, the conditional distribution of the channel admits simple expressions for the first and second moments, summarized below (the proof is omitted due to the space limitation and will be provided elsewhere).
@@
When the phase error $\varepsilon_{k,l}\sim\mathcal{U}[-\delta,\delta]$ vanishes $\delta\to 0$, corresponding to perfect phase tracking, the factor $\rho_{\mathrm{error}}$ approaches one. Conversely, when the phase error is maximally uncertain $\delta\to\pi$, reflecting completely unknown phases, the factor $\rho_{\mathrm{error}}$ tends to zero. In general, imperfect tracking rotates the deterministic LoS component $\bm{\bar{h}}_{k,l}$ by $\hat{\theta}_{k,l}$, and it reduces its magnitude by the factor $\rho_{\mathrm{error}}$, which we refer to as the phase noise penalty.
@@ III Channel Estimation
For random vectors $\bm{x}$ and $\bm{y}$ and a conditioning random quantity $\bm{z}$ (e.g., phase estimates $\hat{\bm{\theta}}$), the conditionally linear MMSE estimator of $\bm{x}$, given $\bm{y}$ and $\bm{z}$, is the affine function of $\bm{y}$ that best approximates $\bm{x}$ in the mean-squared sense, conditioned on $\bm{z}$. It is given by
$\hat{\bm{x}}(\bm{y}\mid\bm{z})=\mathbb{E}[\bm{x}\mid\bm{z}]+\bm{R}_{xy|\bm{F}}\bm{R}_{yy|\bm{z}}^{-1}\big(\bm{y}-\mathbb{E}[\bm{y}\mid\bm{z}]\big),$
@@
Using the measured signal $\bm{y}_{t_{k},l}^{\mathrm{pilot}}$ and conditioning on the phase estimates $\bm{\hat{\theta}}_{l}$, the conditionally linear MMSE estimator of $\bm{h}_{k,l}$ is given by (the proof is omitted due to space limitations and will be provided elsewhere) $(\forall k\in\mathcal{K})(\forall l\in\mathcal{L})$
@@
where
$\bm{\Psi}_{t_{k},l}$ $=\left(\sum_{i\in\mathcal{P}_{k}}\eta_{i}\tau_{p}\bm{\Sigma}_{i,l}+\sigma^{2}\bm{I}_{N}\right)^{-1},$
@@
and $\bm{\Sigma}_{k,l}$ is given in Lemma 1. Furthermore, the estimation error $\bm{\xi}_{k,l}=\bm{{h}}_{k,l}-\bm{\hat{h}}_{k,l}$ has zero mean and the covariance matrix given by
@@
We assume that all statistical parameters, such as $\bm{\bar{h}}_{k,l}$, $\bm{R}_{k,l}$, $\sigma^{2}$ and $\rho_{\mathrm{error}}$ are known across the network. The resulting estimator $\bm{\hat{h}}_{k,l}$ is affine in the received signal $\bm{y}_{t_{k},l}$, but nonlinear in the phase estimate $\bm{\hat{\theta}}_{l}$.
@@ IV-F Ergodic Rates
Under imperfect phase tracking, both the estimated phase $\hat{\theta}_{k,l}$ and the true phase $\theta_{k,l}$ are random quantities that are statistically related through the phase estimation error. We evaluate the network performance using two widely used bounds on the achievable spectral efficiency (SE) of user kk—namely, the lower bound based on use-and-then-forget (UatF) [3] decoding and the upper bound based on the optimistic ergodic rate (OER), introduced as a genie-aided bound in [3]. These bounds remain valid under imperfect phase tracking, since the expectations in these bounds implicitly average over all sources of randomness, including small-scale fading and LoS phase estimates.
@@
The UatF bound models the case where CSI is exploited for beamforming but not for decoding. The ergodic SE $R_{k}$ of user $k$ is lower bounded by
@@
with effective SINR
@@
where $\bm{g}_{ki}=\sqrt{p_{i}}\bm{v}^{\mathsf{H}}_{k}\bm{D}_{k}\bm{h}_{i}$.
@@
This provides an upper bound on the achievable SE $R_{k}$ and a useful benchmark for assessing the tightness of lower bounds. It is given by
@@ V Numerical Results
Figure 1: Comparison of UL SEs achieved for centralized and distributed beamforming under different LOS phase tracking scenarios. Fig 1a) CDF of per-user SE in centralized cell-free using OER bound (16). Fig 1b) CDF of per-user SE in distributed cell-free using OER bound (16). In both (a) and (b), the black line refers to two cases from literature, the optimistic case where the phase is perfectly known (Cor. 1), the pessimistic case where the phase is completely unknown (Cor. 2), and the dotted lines represent our proposed model in Proposition 1 where the LoS phase is partially trackable with some error uniformly distributed as $\sim\mathcal{U}[-\delta,\delta]$. We compare for $\delta\in[0^{\circ},15^{\circ},30^{\circ},45^{\circ},180^{\circ}]$ with all the channels $\kappa=5$. Fig 1c) shows the average SE of the network versus phase uncertainty for centralized and distributed beamformers at different Rician factors $\kappa$ using UatF bound (15). We plot for $\kappa\in[1,5,20,100]$ and $\delta\in[0^{\circ},15^{\circ},30^{\circ},45^{\circ},90^{\circ},180^{\circ}]$.
@@ V-A Parameters and Setup
Network area $d$ × $d$, 10001000 m x $1000$ m
@@
Maximum UL transmit power $p_{\mathrm{max}}=100$ mW
@@
Shadow fading LoS $\sigma_{\mathrm{sf}}=8$ dB
@@
$\beta_{k,l}=35.4-20\log_{10}(f_{c})-26\log_{10}\!\left(\tfrac{d_{k,l}}{1\text{m}}\right)+F_{k,l},[\text{dB}],$ (17)
@@ V-B Results and Conclusions
Figures 1a and 1b show the cumulative distribution functions (CDFs) of the per-user uplink SE, computed using the OER bound (16), for centralized (Section IV-D) and distributed (Section IV-E) beamforming in a cell-free network, respectively. We consider LoS propagation and fix the Rician factor at $\kappa=5$ for all channels. The black curves represent the two benchmarks from the literature: the optimistic case of perfectly known LoS phases (Corollary 1) and the pessimistic case of completely unknown phases (Corollary 2). The intermediate cases, corresponding to imperfect phase tracking captured by our proposed model in Proposition 1, are illustrated by dotted lines for $\delta\in[0^{\circ},15^{\circ},30^{\circ},45^{\circ},180^{\circ}]$, where the phase error $\varepsilon\sim\mathcal{U}[-\delta,\delta]$. The results demonstrate that even partial knowledge of the slowly varying LoS phases provides substantial performance gains compared to the no-phase-knowledge scenario. As $\delta$ increases, the gap to the perfect knowledge benchmark widens; however, moderate errors (e.g., $\delta=15^{\circ}$) already yield distributions close to the ideal case for both architectures. Noticeable degradation appears only for large phase uncertainties, particularly in the distributed setting.
@@
Figure 1c illustrates the average network SE as a function of the phase uncertainty, evaluated using the UatF bound (15). The results compare centralized and distributed beamforming under different Rician factors $\kappa\in[1,5,20,100]$. At high Rician factors (e.g., $\kappa=100$), both designs tend to converge, and the performance gap between centralized and distributed cell-free network becomes negligible when the LoS phases are accurately known. This observation is consistent with our previous findings [2], where we showed that under strong LoS conditions and perfect phase knowledge, distributed beamforming that exploits the statistical channel information can closely approach the centralized benchmark. In the presence of phase uncertainty, however, the achievable rates of both architectures degrade, with the impact becoming more pronounced as $\kappa$ increases. Nevertheless, even coarse or partially accurate phase tracking provides significant performance gains over the no-phase knowledge model, particularly for distributed designs.
@@
Overall, the results demonstrate that partial knowledge of LoS phases remains highly beneficial for both centralized and distributed beamforming, with centralized more robust under severe phase errors. These findings underscore the importance of phase-aware channel estimation and beamforming strategies for practical 6G cell-free massive MIMO deployments.

### v2

@@ Abstract
We study the impact of imperfect line-of-sight (LoS) phase tracking on the uplink performance of cell-free massive MIMO networks. Unlike prior works that assume perfectly known or completely unknown phases, we consider a realistic regime where LoS phases are estimated with residual uncertainty due to hardware impairments, mobility, and synchronization errors. To this end, we propose a Rician fading model where LoS components are rotated by imperfect phase estimates and attenuated by a deterministic phase-error penalty factor.
We derive a linear MMSE channel estimator that accounts for statistical phase errors and unifies prior results, reducing to the Bayesian MMSE estimator when phase is perfectly known and to a zero-mean model when no phase information is available. To address the non-Gaussian setting, we introduce a virtual uplink model that preserves second-order statistics of channel estimation, enabling the derivation of tractable virtual centralized and distributed MMSE beamformers. To ensure fair assessment of network performance, we apply these virtual beamformers to the operational uplink model that reflects the actual physical channel and compute the spectral efficiency bounds available in the literature.
@@ I Introduction
Existing studies typically analyze cell-free systems under two limiting cases of LoS phase knowledge. On one hand, studies such as [17, 2] and phase-aware case in [12, 16] assume perfect knowledge of the LoS phases, which enables tractable analysis but overlooks practical estimation errors. On the other hand, studies [16, 12] also consider the case of completely unknown LoS phases, thereby discarding valuable deterministic information inherent in the slowly varying LoS components. However, as also pointed out in [16, 12], the practical system performance fall in between these two scenarios. Furthermore, the performance evaluation of LoS-aware cell-free systems has largely been restricted to maximum-ratio (MR) combining [17, 16, 12]. While analytically convenient, MR combining does not fully exploit the beamforming potential of large-scale cell-free networks. Complementary studies [13, 5, 14] have investigated phase noise effects in Rayleigh fading channels, that is, without explicit modeling of LoS components. These simplifications limit our understanding of the impact of partial LoS phase knowledge on channel estimation and beamforming in cell-free systems.
@@ II System and Channel Model
We consider a user-centric cell-free massive multiple-input multiple-output (mMIMO) network with $L$ access points, indexed by $\mathcal{L}:=\{1,\dots,L\}$, jointly serving $K$ users indexed by $\mathcal{K}:=\{1,\dots,K\}$. The positions are fixed and arbitrarily distributed across the service area. All APs are connected to a central processing unit (CPU) via an error-free fronthaul network [10]. Each AP is equipped with $N$ antennas, while each user has a single antenna.
@@ II-A Rician Channel under Imperfect Phase Tracking
In an arbitrary coherence block, let $\bm{h}_{k,l}\in\mathbb{C}^{N}$ denote a realization of the channel vector between the user $k\in\mathcal{K}$ and the $N$ antennas of AP $l\in\mathcal{L}$. More specifically, the channel is modeled as $(\forall k\in\mathcal{K})(\forall l\in\mathcal{L})$
@@
where $\bm{\bar{h}}_{k,l}\in\mathbb{C}^{N}$ is the deterministic spatial signature of the LoS component, and $\mathbf{R}_{k,l}\in\mathbb{C}^{N\times N}$ is the spatial covariance matrix of the non-line-of-sight (NLoS) components $\bm{\tilde{h}}_{k,l}$, and $\theta_{k,l}\sim\mathcal{U}[-\pi,\pi]$ is the phase of the LoS component. For each realization of $\theta_{k,l}$, an estimate
$\hat{\theta}_{k,l}=\theta_{k,l}-\varepsilon_{k,l}$ (2)
@@
However, as noted in [12, 16], practical systems operate between these extremes. In fact, the LoS phase often varies more slowly than the NLoS components, whose dynamics are dominated by delay and Doppler spreads. For example, recent work [17] claims that the LoS phase may remain approximately constant across multiple coherence blocks, particularly in the frequency domain. Therefore, with our model we focus on the intermediate case where the LoS phase $\theta_{k,l}$ varies more slowly than the NLoS components $\bm{\tilde{h}}_{k,l}$, but still fast enough such that it cannot be tracked with desired accuracy. We denote by $\bm{\hat{\theta}}_{l}=[\hat{\theta}_{1,l},\dots,\hat{\theta}_{K,l}]^{\mathsf{T}}\in\mathbb{C}^{1\times K}$ and $\bm{\hat{\theta}}=[\bm{\hat{\theta}}_{1},\dots,\bm{\hat{\theta}}_{L}]\in\mathbb{C}^{L\times K}$ the phase estimates for all channels of AP $l$ and the network respectively.
The phase estimates $\bm{\hat{\theta}}$ are assumed to be available as common knowledge across the network. The phase estimation error $\varepsilon_{k,l}$ in (2) is modeled as uniformly distributed, $\varepsilon_{k,l}\sim\mathcal{U}[-\delta,\delta]$, with $0\leq\delta\leq\pi$, and it is assumed independent of the NLoS component. Under these assumptions, the conditional distribution of the channel admits simple expressions for the first and second moments, summarized below (the proof is provided in the Appendix VI-A).
@@
When the phase error $\varepsilon_{k,l}\sim\mathcal{U}[-\delta,\delta]$ vanishes as $\delta\to 0$, corresponding to perfect phase tracking, the factor $\rho_{\mathrm{error}}$ approaches one. Conversely, when the phase error becomes maximally uncertain as $\delta\to\pi$, reflecting completely unknown phases, the factor $\rho_{\mathrm{error}}$ tends to zero. In general, imperfect tracking rotates the deterministic LoS component $\bm{\bar{h}}_{k,l}$ by $\hat{\theta}_{k,l}$, and it reduces its magnitude by the factor $\rho_{\mathrm{error}}$, which we refer to as the phase noise penalty.
@@ III Channel Estimation
For random vectors $\bm{x}$ and $\bm{y}$, and a conditioning random quantity $\bm{z}$ (e.g., phase estimates $\hat{\bm{\theta}}$), the conditionally linear MMSE estimator of $\bm{x}$, given $\bm{y}$ and $\bm{z}$, is the affine function of $\bm{y}$ that minimizes the conditional mean-squared error $\mathbb{E}\big[\lVert\bm{x}-\bm{\hat{x}}\rVert^{2}\!\mid\bm{z}\big]$. It is given by
$\hat{\bm{x}}(\bm{y}\mid\bm{z})=\mathbb{E}[\bm{x}\mid\bm{z}]+\bm{R}_{xy|\bm{z}}\bm{R}_{yy|\bm{z}}^{-1}\big(\bm{y}-\mathbb{E}[\bm{y}\mid\bm{z}]\big),$
@@
Using the measured signal $\bm{y}_{t_{k},l}^{\mathrm{pilot}}$ and conditioning on the phase estimates $\bm{\hat{\theta}}_{l}$, the conditionally linear MMSE estimator of $\bm{h}_{k,l}$ is given by (the proof is provided in the Appendix VI-B) $(\forall k\in\mathcal{K})(\forall l\in\mathcal{L})$
@@
$\text{where}\quad\bm{\Psi}_{t_{k},l}$ $=\left(\sum_{i\in\mathcal{P}_{k}}\eta_{i}\tau_{p}\bm{\Sigma}_{i,l}+\sigma^{2}\bm{I}_{N}\right)^{-1},$
@@
and $\bm{\Sigma}_{k,l}$ is the conditional covariance matrix of channel $\bm{h}_{k,l}$, given in Lemma 1. Furthermore, the estimation error $\bm{\xi}_{k,l}=\bm{{h}}_{k,l}-\bm{\hat{h}}_{k,l}$ has zero mean and the covariance matrix given by
@@
We assume that all statistical parameters such as $\bm{\bar{h}}_{k,l}$, $\bm{R}_{k,l}$, $\sigma^{2}$ and $\rho_{\mathrm{error}}$ are known across the network. The resulting estimator $\bm{\hat{h}}_{k,l}$ is affine in the received signal $\bm{y}_{t_{k},l}$, but nonlinear in the phase estimate $\bm{\hat{\theta}}_{l}$.
@@ IV-F Ergodic Rates
Under imperfect phase tracking, both the estimated phase $\hat{\theta}_{k,l}$ and the true phase $\theta_{k,l}$ are random quantities that are statistically related through the phase estimation error. To provide a comprehensive performance assessment, we evaluate the system using two widely used bounds on the achievable spectral efficiency (SE)- namely the lower bound based on use-and-then-forget (UatF) [3] decoding and the upper bound based on the optimistic ergodic rate (OER), introduced as a genie-aided bound in [3]. These bounds remain valid under imperfect phase tracking, since the expectations in these bounds implicitly average over all sources of randomness, including small-scale fading and LoS phase estimates.
@@
The UatF bound models the case where instantaneous CSI is exploited for beamforming but not for decoding. The ergodic SE $R_{k}$ of user $k$ is lower bounded by [3]
@@
where the effective SINR is given by
@@
and $\bm{g}_{ki}=\sqrt{p_{i}}\bm{v}^{\mathsf{H}}_{k}\bm{D}_{k}\bm{h}_{i}$ is the effective channel.
@@
To benchmark the performance limits, we consider the OER bound, which assumes a ‘genie-aided’ decoder that has perfect knowledge of the instantaneous effective channel (including the realization of the phase error), while treating interference and noise as uncorrelated Gaussian noise. The ergodic SE $R_{k}$ of user $k$ is upper bounded by [3]
@@ V Numerical Results
Figure 1: Comparison of uplink SE obtained using the OER bound (17) under different LoS phase-tracking scenarios. Figure 1a shows the CDF of the per-user SE for centralized cell-free beamforming, while Figure 1b shows the corresponding CDF for distributed cell-free beamforming. In both (a) and (b), the black solid curves represent two benchmark cases from the literature: the optimistic case with perfectly known LoS phases (Cor. 1) and the pessimistic case with completely unknown phases (Cor. 2). The dotted curves correspond to the proposed model in Proposition 1, where the LoS phase is partially tracked with an error $\varepsilon\sim\mathcal{U}[-\delta,\delta]$. Results are shown for $\delta\in\{0^{\circ},15^{\circ},30^{\circ},45^{\circ},180^{\circ}\}$ with a fixed Rician factor $\kappa=5$ for all channels. Figure 1c shows the average network SE versus phase uncertainty for centralized and distributed beamforming at different Rician factors $\kappa\in\{1,5,20,100\}$ and $\delta\in\{0^{\circ},15^{\circ},30^{\circ},45^{\circ},90^{\circ},180^{\circ}\}$.
Figure 2: Comparison of upper (OER) and lower (UatF) bounds on the achievable SE under imperfect LoS phase tracking for centralized beamforming in Figure 2a and distributed beamforming in Figure 2b. In both (a) and (b), solid curves correspond to the OER bound, while dotted curves represent the UatF bound. Results are shown for $\delta\in\{0^{\circ},45^{\circ},180^{\circ}\}$ with a fixed Rician factor $\kappa=5$ for all channels. Figure 2c shows the average network SE versus phase uncertainty for centralized and distributed beamforming at Rician factors $\kappa\in\{1,20\}$, using the OER and UatF bounds (solid and dotted curves, respectively).
@@ V-A Parameters and Setup
Network area $1000\,\text{m}\,\times 1000\,\text{m}$
@@
Maximum uplink transmit power $p_{\mathrm{max}}=100$ mW
@@
Shadow fading deviation $\sigma_{\mathrm{sf}}=8$ dB
@@
$\beta_{k,l}=35.4-20\log_{10}(f_{c})-26\log_{10}\!\left(\tfrac{d_{k,l}}{1\text{m}}\right)+F_{k,l},\quad[\text{dB}],$ (18)
@@ V-B Results and Conclusions
Figures 1a and 1b show the cumulative distribution functions (CDFs) of the uplink SE per-user, computed using the OER bound (17), for centralized (Section IV-D) and distributed (Section IV-E) beamforming in a cell-free network, respectively. We consider LoS propagation and fix the Rician factor at $\kappa=5$ for all channels. The black curves represent the two benchmarks from the literature: the optimistic case of perfectly known LoS phases (Corollary 1) and the pessimistic case of completely unknown phases (Corollary 2). The intermediate cases, corresponding to the imperfect phase tracking captured by our proposed model in Proposition 1, are illustrated by dotted curves for $\delta\in\{0^{\circ},15^{\circ},30^{\circ},45^{\circ},180^{\circ}\}$, where the phase error is $\varepsilon\sim\mathcal{U}[-\delta,\delta]$. The results demonstrate that even partial knowledge of the slowly varying LoS phases provides substantial performance gains compared to the no-phase-knowledge scenario. As $\delta$ increases, the gap to the perfect knowledge benchmark widens; however, moderate errors (e.g., $\delta=15^{\circ}$) already result in distributions close to the ideal case for both architectures. Noticeable degradation appears only for large phase uncertainties, particularly in a distributed setting.
@@
Figure 1c illustrates the average network SE as a function of the phase uncertainty, evaluated using the OER bound (17). The results compare centralized and distributed beamforming under different Rician factors $\kappa\in\{1,5,20,100\}$. At high Rician factors (e.g., $\kappa=100$), both designs tend to converge, and the performance gap between centralized and distributed cell-free networks becomes negligible when the LoS phases are accurately known. This observation is consistent with our previous findings [2], where we showed that under strong LoS conditions and perfect phase knowledge, distributed beamforming that exploits the statistical channel information can closely approach the centralized benchmark. However, in the presence of phase uncertainty, the achievable rates of both architectures degrade, with the impact becoming more pronounced as $\kappa$ increases. Nevertheless, even coarse or partially accurate phase tracking provides significant performance gains over the no-phase knowledge model, particularly for distributed designs.
In Figure 2, we evaluate both $R_{k}^{\text{uatf}}$ (16) and $R_{k}^{\text{oer}}$ (17) under imperfect LoS phase tracking to assess the tightness of the SE bounds. The gap between the two bounds is observed to be small, particularly for $\kappa=20$ in Figure 2c with $\delta\in\{0^{\circ},15^{\circ}\}$. This behavior indicates that due to the dominant LoS component with an almost perfectly known LoS phase, the effective channel $\bm{g}_{kk}$ behaves nearly deterministically, meaning that the lack of instantaneous CSI at the decoder results in a negligible performance loss.
@@
Overall, the results demonstrate that partial knowledge of LoS phases remains highly beneficial for both centralized and distributed beamforming, with centralized beamforming exhibiting greater robustness under severe phase errors. These findings underscore the importance of phase-aware channel estimation and beamforming strategies for practical 6G cell-free massive MIMO deployments.
VI Appendix
This appendix provides detailed proofs of Lemma 1 and Proposition 1, which were omitted from the main text.
VI-A Proof of Lemma 1
We derive the first and second moments of the channel for a specific AP-user pair. For brevity, we drop the user and AP indices $\{k,l\}$ and denote the phase error by $\epsilon\sim\mathcal{U}[-\delta,\delta]$. The characteristic function of the uniform phase error is calculated as
$\rho_{\text{error}}\triangleq\mathbb{E}[e^{j\epsilon}]$ $=\frac{1}{2\delta}\int_{-\delta}^{\delta}e^{jx}dx=\begin{cases}1,&\delta=0,\\ \frac{\sin(\delta)}{\delta},&0<\delta\leq\pi.\end{cases}$
Note that for the symmetric interval $\delta,\delta]$, $\rho_{\text{error}}$ is real-valued.
1) Conditional Expectation: Substituting $\theta=\hat{\theta}+\epsilon$ into the channel model (1), we have $\mathbf{h}=\bar{\mathbf{h}}e^{j(\hat{\theta}+\epsilon)}+\tilde{\mathbf{h}}$. The terms $\bar{\mathbf{h}}$ and $e^{j\hat{\theta}}$ (conditioned on the estimate $\hat{\theta}$) are deterministic, while $\tilde{\mathbf{h}}$ remains zero-mean and independent of the phase error. Taking the conditional expectation yields
$\bm{\mu}\triangleq\mathbb{E}[\mathbf{h}\mid\hat{\theta}$] $=\bar{\mathbf{h}}e^{j\hat{\theta}}\mathbb{E}[e^{j\epsilon}]+\mathbb{E}[\tilde{\mathbf{h}}]=\bar{\mathbf{h}}e^{j\hat{\theta}}\rho_{\text{error}}.$
2) Conditional Covariance: The conditional covariance matrix is defined as
$\bm{\Sigma}=\mathbb{E}[(\mathbf{h}-\bm{\mu})(\mathbf{h}-\bm{\mu})^{\mathsf{H}}\mid\hat{\theta}$], (19)
where the deviation from the expectation is given by $\mathbf{h}-\bm{\mu}=\bar{\mathbf{h}}e^{j\hat{\theta}}(e^{j\epsilon}-\rho_{\text{error}})+\tilde{\mathbf{h}}$. Since $\epsilon$ and $\tilde{\mathbf{h}}$ are independent and $\tilde{\mathbf{h}}$ is zero-mean, the cross-terms in the expectation vanish. Equation (19) simplifies to
$\bm{\Sigma}$ $=\bar{\mathbf{h}}\bar{\mathbf{h}}^{\mathsf{H}}\mathbb{E}[|e^{j\epsilon}-\rho_{\text{error}}|^{2}]+\mathbb{E}[\tilde{\mathbf{h}}\tilde{\mathbf{h}}^{\mathsf{H}}]$
$=\bar{\mathbf{h}}\bar{\mathbf{h}}^{\mathsf{H}}(1-\rho_{\text{error}}^{2})+\mathbf{R}\qquad\text{(using }|e^{j\epsilon}|^{2}=1\text{)},$ (20)
which concludes the proof. $\blacksquare$
VI-B Proof of Proposition 1
We derive the conditionally linear MMSE estimator for a specific AP-user pair. For clarity, we omit the user index $k$, AP index $l$, and pilot index $t_{k}$. Let $\mathbf{h}$ denote the channel of the user of interest, $\hat{\theta}$ the phase estimate, and $\mathbf{y}$ the observed signal for the assigned pilot.
1) Preliminaries: From (4), the received pilot signal from the set of pilot-sharing users $\mathcal{P}$ is given by $\mathbf{y}=\sum_{i\in\mathcal{P}}\sqrt{\eta_{i}}\tau\mathbf{h}_{i}+\mathbf{n},$ where $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\tau\mathbf{I}_{N})$. Using the result $\mathbb{E}[\mathbf{h}_{i}|\hat{\theta}_{i}]=\bm{\mu}_{i}$ from the proof of Lemma 1, the conditional expectation of the observed signal is
$\bar{\mathbf{y}}\triangleq\mathbb{E}[\mathbf{y}|\hat{\bm{\theta}}]=\sum_{i\in\mathcal{P}}\sqrt{\eta_{i}}\tau\bm{\mu}_{i}.$ (21)
The conditional cross-covariance between the target channel $\mathbf{h}$ and the observation $\mathbf{y}$ is $\mathbf{C}_{hy}=\mathbb{E}[(\mathbf{h}-\bm{\mu})(\mathbf{y}-\bar{\mathbf{y}})^{\mathsf{H}}\mid\hat{\bm{\theta}}$]. Exploiting the independence of user channels and receiver noise, it simplifies to
$\mathbf{C}_{hy}=\mathbb{E}[(\mathbf{h}-\bm{\mu})(\sqrt{\eta}\tau(\mathbf{h}-\bm{\mu}))^{\mathsf{H}}\mid\hat{\bm{\theta}}]=\sqrt{\eta}\tau\bm{\Sigma},$ (22)
where $\eta$ and $\bm{\Sigma}$ refer to the parameters of the user of interest.
Similarly, exploiting the independence of user channels and receiver noise, the conditional covariance matrix of the observed signal is derived as
$\mathbf{C}_{yy}=\mathbb{E}[(\mathbf{y}-\bar{\mathbf{y}})(\mathbf{y}-\bar{\mathbf{y}})^{\mathsf{H}}\mid\hat{\bm{\theta}}]=\sum_{i\in\mathcal{P}}\eta_{i}\tau^{2}\bm{\Sigma}_{i}+\sigma^{2}\tau\mathbf{I}_{N},$ (23)
where $\bm{\Sigma}_{i}$ is the conditional channel covariance of user $i$ defined in Lemma 1.
2) Estimator Derivation: Substituting the moments derived above into the Definition (1) of conditionally linear MMSE estimator gives
$\hat{\mathbf{h}}=\bm{\mu}+\mathbf{C}_{hy}\mathbf{C}_{yy}^{-1}(\mathbf{y}-\bar{\mathbf{y}}).$ (24)
To recover the expression in Proposition 1, $\tau$ can be factored out as follows
$\mathbf{C}_{hy}\mathbf{C}_{yy}^{-1}$ $=(\sqrt{\eta}\tau\bm{\Sigma})\big(\tau\big[\sum_{i\in\mathcal{P}}\eta_{i}\tau\bm{\Sigma}_{i}+\sigma^{2}\mathbf{I}_{N}\big]\big)^{-1}$
$=\sqrt{\eta}\bm{\Sigma}\underbrace{\big(\sum_{i\in\mathcal{P}}\eta_{i}\tau\bm{\Sigma}_{i}+\sigma^{2}\mathbf{I}_{N}\big)^{-1}}_{\triangleq\bm{\Psi}}.$ (25)
$\text{Thus,}\qquad\hat{\mathbf{h}}=\bm{\mu}+\sqrt{\eta}\bm{\Sigma}\bm{\Psi}(\mathbf{y}-\bar{\mathbf{y}}).$ (26)
Finally, the error covariance matrix for a linear MMSE estimator is given by $\mathbf{C}=\bm{\Sigma}-\mathbf{C}_{hy}\mathbf{C}_{yy}^{-1}\mathbf{C}_{hy}^{\mathsf{H}}$ [6], which simplifies to
$\mathbf{C}=\bm{\Sigma}-\eta\tau\bm{\Sigma}\bm{\Psi}\bm{\Sigma}.$ (27)
This concludes the proof. $\blacksquare$